Let a sample space be partitioned into three mutually exclusive and exhaustive events, B1, B2, and, B3. Complete the following probability table. (Round your answers to 2 decimal places.)

Prior
Probabilities
Conditional
Probabilities
Joint
Probabilities
Posterior
Probabilities
P(B1) = 0.11 P(A | B1) = 0.45 P(AB1) = P(B1 | A) =
P(B2) = P(A | B2) = 0.62 P(AB2) = P(B2 |A) =
P(B3) = 0.38 P(A | B3) = 0.85 P(AB3) = P(B3 | A) =
Total = P(A) = Total =
Answer #1

Answer:

Given,

P(B2) = 1 - P(B1) - P(B3)

= 1 - 0.11 - 0.38

= 0.51

Total = 1

P(A \cap B1) = P(B1)*P(A|B1) = 0.11*0.45 = 0.0495

P(A \cap B2) = P(B2)*P(A|B2) = 0.51*0.62 = 0.3162

P(A \cap B3) = P(B3)*P(A|B3) = 0.38*0.85 = 0.323

P(A) = 0.0495 + 0.3162 + 0.323 = 0.6887

P(B1|A) = P(A \cap B1)/P(A) = 0.0495/0.6887 = 0.0719

P(B2|A) = P(A \cap B2)/P(A) = 0.3162/0.6887 = 0.4591

P(B3|A) = P(A \cap B3)/P(A) = 0.323/0.6887 = 0.4690

Total = 1